OpenStudy (anonymous):

find a point slope equation of the line having the given slope and containing the given point. m=-9,(8,9) the equation of the line in a point - slope form is y- = (x- ) the spaces are blanks in the equation

OpenStudy (anonymous):

(y-y1)=m(x-x1) (y-9)=-9(x-8) y=-9x--72+9 y=-9x+81 -9x-y+81=0

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np.